(1/cosx)-(cosx/1+sinx)=tanx
Answer by Alan3354(21605)

(1/cosx)-(cosx/1+sinx)=tanx
1 - (cos^2/(1+sin) = sin
1+sin - cos^2 = sin + sin^2
1+sin = sin + sin^2+cos^2
1+sin = 1+sin